OpenStudy (anonymous):

x^2-12x+30 using quadratic equation

OpenStudy (anonymous):

have to find the value of x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use the formula,in equation \[ax^2+bx+c=0 \] the value of x\[x=-b \pm \sqrt{b^2-4ac}/2a\]

OpenStudy (anonymous):

I did do I leave it at 6+/-sqare root of 12

OpenStudy (anonymous):

\[\sqrt{12}=2\sqrt{3}\]

OpenStudy (anonymous):

oh sorry,